OpenStudy (anonymous):

x^2-14x+4=0 solve for x

OpenStudy (amistre64):

(x^2 -14x +7^2)+4-7^2=0 (x-7)^2 = 49 - 4 x-7 = +- sqrt(45) x = 7 +- 3sqrt(5)